# Hyperparameters are all you need: Using five-step inference for an original diffusion model to generate images comparable to the latest distillation model

### v1

@@ 4.1 Performance
In this study, we employ $p_{1}=7$, $p_{2}=1.2$, $stop=3$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images. For the eight-step inference, as the precise discretization method, we can completely avoid using free-U, which can enhance the aesthetic of images while reducing its FID in the few-step inference. We report the results of the eight-step inference in the COCO 2014, COCO 2017, and LAION dataset in Tab. 1, Tab. 2, and Tab. 3. Moreover, we use $p_{1}=5$, $p_{2}=1.2$ as the setting in the five- and six-step inference of generating 512 x 512 images and the six-step inference of generating 1024 x 1024 images. The result of the five- and six-step inference that synthesizes 512 x 512 images is also given in Tab. 1, Tab. 2, and Tab. 3. Moreover, the result of the 1024 x 1024 images is given in Tab. 4, Tab. 5, and Tab. 6. The free-U is used throughout the inference when we sample a 512 x 512 image, and it is used in the last three steps when we synthesize a 1024 x 1024 image. Moreover, the ODE solver, based on the analysis in Section 3, is DPM++2m for eight-step inference and DPM++1s for five-step inference. The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively. In the generation of 1024 x 1024 images in 6 steps, we use the SDE-Editing method mentioned in the paper[27]. Furthermore, all the synthesized images in Tab. 1 and Tab. 4 employs the random 10k COCO caption as a guidance prompt. In Tab. 2 and Tab. 5, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance. Furthermore, we use 10k captions in LAION for Tab. 3 and Tab. 6. We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2[43, 3, 19, 32]. Our approach outperforms state-of-the-art models for most of the datasets. However, some of the models[43] are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison. The $w$ in the Tab. 1 and Tab. 4 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
@@
Original DPM++2m(20) 17.3 30.97
@@
custom DPM++1s (6) $w=7.5$ 23 31.23
custom DPM++1s (6) $w=5.5$ 24.25 31.49
@@
custom DPM++1s (6) $w=7.5$ 29.52 31.15
custom DPM++1s (6) $w=5.5$ 30.51 31.47
@@
custom DPM++1s (6) $w=7.5$ 20.94 31.01
custom DPM++1s (6) $w=5.5$ 22.10 31.59
@@ 5.1 Limitation and the future work
A diffusion model is a robust generative model[4] that can prevent adversarial attacks. Most attacks that hack the latent diffusion model attack its VAE decoder. Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that mainly depends on the property of the VAE. Moreover, in future research, we explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models.
Acknowledgments
We thank LetPub (www.letpub.com.cn) for its linguistic assistance during the preparation of this manuscript.

### v2

@@ 4 Experiments
4.1 Ablation Study
We conduct an ablation Study on a customized UniPC solver that aims to generate a 1024 x 1024 image in a few-step inference guided by 5000 prompts sampling from the COCO 2017 dataset. Explicitly, for the 8-step inference, we employ $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ as the discrete setting. And for the 6-step inference, the discrete setting is $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Then, we try the following setting: 1. apply Free-U on every step of inference, 2. apply Free-U only on the last few steps of inference, 3. Don’t apply Free-U. The experimental result is in the table˜1. The $w$ in the table˜1 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. Meanwhile, for the 8-step UniPC solver, we use a 2nd-order multistep algorithm, whereas the 6-step solver is 1st-order solver.
Solver FID CLIP-Score
custom UniPC(6) apply Free-U on the every steps $w=5.5$ 28.88 30.96
custom UniPC(8) apply Free-U on the every steps $w=5.5$ 33.70 30.50
custom UniPC(6) don’t apply Free-U $w=5.5$ 27.13 31.49
custom UniPC(8) don’t apply Free-U $w=5.5$ 25.47 31.39
custom UniPC(6) start apply Free-U at 4th step $w=5.5$ 25.20 31.42
custom UniPC(8) start apply Free-U at 6th step $w=5.5$ 24.38 31.38
custom UniPC(6) apply Free-U on the every steps $w=7.5$ 39.62 30.33
custom UniPC(8) apply Free-U on the every steps $w=7.5$ 54.26 29.63
custom UniPC(6) don’t apply Free-U $w=7.5$ 25.34 31.53
custom UniPC(8) don’t apply Free-U $w=7.5$ 24.61 31.22
custom UniPC(6) start apply Free-U at 4th step $w=7.5$ 25.29 31.44
custom UniPC(8) start apply Free-U at 6th step $w=7.5$ 24.58 31.21
Table 1: Results of applying Free-U in different steps of inference.
@@ 4.2 Performance
In this study, we employ our algorithm on the DPM++ 2m solver and UniPC solver. We employ $p_{1}=7$, $p_{2}=1.2$, $stop=8$, $M=12$, and $N=8$ as the hyperparameters of the discrete setting in the eight-step inference when generating 1024 x 1024 and 512 x 512 images via DPM++ 2m. Additionally, $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=8$ are applied on the custom UniPC solver in the eight-step inference when generating 1024 x 1024 images. And for the six-step inference for generating 1024 x 1024-resolution images, the hyperparameters of the UniPC solver are $p_{1}=12$, $p_{2}=1.2$, $stop=10$, $M=12$, and $N=6$. Meanwhile, for 512 x 512 images, the hyperparameters of the 8-step UniPC solver are $p_{1}=7$, $p_{2}=1.2$, $stop=9$, $M=12$, and $N=8$. And for the customized 5-step UniPC solver for 512 x 512 images, the hyperparameters are $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=5$. Then, the 6-step UniPC solver of the 512 x 512-resolution images has hyperparameters $p_{1}=5$, $p_{2}=1.2$, $stop=6$, $M=8$, and $N=6$.
For the eight-step inference that uses DPM++ 2m to generate 512x512-resolution images, we avoid using free-U. And for the customized DPM++ 2m solver that generates 1024x1024-resolution images, the $t_{aug}$ apply free-U is 8. For all UniPC solver customized to generate 512x512 resolution images, $t_{aug}=4$. And $t_{aug}=6$, $t_{aug}=4$ for 6-step UniPC solver and 8-step UniPC solver customeized for 1024 x 1024-resolution images, respectively. And we only use AFS on the UniPC solver customized for the 1024 x 1024-resolution image. We don’t guarantee we provide the best hyperparameters; we just ensure we provide hyperparameters that can sample images with FID performance comparable to that of the inference result from the 20-step DPM++ solver and the SOTA distillation algorithm.
We report the results of the inference that generate 512 x 512-resolution images in the COCO 2014, COCO 2017, and LAION dataset in table˜2, table˜3, and table˜4. Moreover, the result of the 1024 x 1024 images is given in table˜5, table˜6, and table˜7. Explicitly, the UniPC solver is a 2nd-order multistep solver for eight-step inference and a 1st-order multistep solver for five and six-step inference. The inference model that uses our algorithm for synthesizing 512 x 512 images and 1024 x 1024 images is Stable Diffusion v1.5 and Dreamshaper XL, respectively. Furthermore, all the synthesized images in table˜2 and table˜5 employs the random 10k COCO caption as a guidance prompt. In table˜3 and table˜6, for each image in the COCO 2017 dataset, we fetch one particular caption of it as guidance. Furthermore, we use 10k captions in LAION for table˜4 and table˜7. We also compare the FID performance and the clip score of AMED-Pulging solver, progressive distillation, SDXL turbo, SDXL-Lightning, Flash DiffusionXL, Flash Diffusion, and DMD2 (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024); Sauer et al. (2024)). Our approach outperforms state-of-the-art models for most of the datasets. However, some of the models (Yin et al. (2024a); Chadebec et al. (2025); Lin et al. (2024)) are trained in LAION, which do not separate the validation and the training sets, and may affect the comparison. The $w$ in the table˜2 and table˜5 is the guidance scale of inference, and the number in the parentheses is the total number of function evaluations (NFEs) of the model. The $b_{1}$, $b_{2}$, $s_{1}$, and $s_{2}$ in the Free-U decorator is 1.1, 1.1, 0.9, 0.2.
@@
(Our) custom UniPC(5) $w=7.5$ 19.35 30.63
(Our) custom UniPC(6) $w=7.5$ 16.41 30.97
(Our) custom UniPC(8) $w=7.5$ 16.00 31.15
(Our) custom UniPC(5) $w=5.5$ 15.19 30.93
(Our) custom UniPC(6) $w=5.5$ 14.72 31.08
(Our) custom UniPC(8) $w=5.5$ 15.26 31.12
Original DPM++2m(20) $w=7.5$ 17.3 30.97
@@
(Our) custom UniPC(5) $w=7.5$ 22.70 30.99
(Our) custom UniPC(6) $w=7.5$ 24.02 31.04
(Our) custom UniPC(8) $w=7.5$ 24.55 30.99
(Our) custom UniPC(5) $w=5.5$ 21.31 30.92
(Our) custom UniPC(6) $w=5.5$ 20.82 31.04
(Our) custom UniPC(8) $w=5.5$ 21.71 31.07
@@
(Our) custom UniPC(5) $w=7.5$ 20.43 30.87
(Our) custom UniPC(6) $w=7.5$ 17.56 31.63
(Our) custom UniPC(8) $w=7.5$ 17.38 32.08
(Our) custom UniPC(5) $w=5.5$ 17.44 31.73
(Our) custom UniPC(6) $w=5.5$ 16.84 32.08
(Our) custom UniPC(8) $w=5.5$ 17.32 32.26
@@
custom UniPC (8) $w=7.5$ 18.18 31.23
custom UniPC (8) $w=5.5$ 17.87 31.40
custom UniPC (6) $w=7.5$ 18.80 31.42
custom UniPC (6) $w=5.5$ 19.07 31.45
@@
custom UniPC (8) $w=7.5$ 24.58 31.21
custom UniPC (8) $w=5.5$ 24.38 31.38
custom UniPC (6) $w=7.5$ 25.29 31.44
custom UniPC (6) $w=5.5$ 25.20 31.42
@@
custom UniPC (8) $w=7.5$ 21.59 31.73
custom UniPC (8) $w=5.5$ 20.41 32.06
custom UniPC (6) $w=7.5$ 22.51 31.82
custom UniPC (6) $w=5.5$ 22.33 31.87
@@ 5.1 Limitation and the future work
Firstly, a diffusion model is a robust generative model (Chen et al. (2024)) that can prevent adversarial attacks. Most attacks that hack the latent diffusion model attack its VAE decoder. Thus, it remains unclear whether the robustness of the diffusion model will be affected when using our algorithm in a scenario that largely depends on the property of the VAE. Moreover, in future research, we will explore whether it is possible to find a better scheduler in a low-step inference process via an automated algorithm, and whether this new time scheduler is available to different latent diffusion models. Last, the hyperparameters of the ACGN model need further exploration.